OpenStudy (anonymous):

y(s)=1/5s^2 - 6/s find y'(s)

OpenStudy (lgbasallote):

is s^2 in the denom? and s too?

OpenStudy (anonymous):

let me rewrite

OpenStudy (asnaseer):

is this your equation:\[y(s)=\frac{1}{5s^2}-\frac{6}{s}\]

OpenStudy (anonymous):

y(s)=(1)/(5s^2) - (6)/(s)

OpenStudy (anonymous):

YES!

OpenStudy (anonymous):

sorry about the confusion everyone

OpenStudy (lgbasallote):

i see...well the constants don't affect the differentiation..just turn the s to negative exponents then power rule. try

OpenStudy (lgbasallote):

you also dont have to LCD it or anything...itll just make it harder

OpenStudy (asnaseer):

yes - Igbasallote is correct. rewrite the equation as:\[y(s)=\frac{1}{5}s^{-2}-6s^{-1}\]you should then be able to use the standard rules of differentiation.

OpenStudy (asnaseer):

i.e. use the rule that if:\[y=ax^b\]then:\[y'=abx^{b-1}\]

OpenStudy (asnaseer):

do you understand?

OpenStudy (anonymous):

YES IT DOES! THANKS!

OpenStudy (asnaseer):

yw

OpenStudy (anonymous):

:D